# Quasinormal modes of nonlocal gravity black holes

### v1

@@ I Introduction
Attempts to address these challenges have led to the development of modified gravity scenarios Ferraro and Fiorini (2007); Sotiriou and Faraoni (2010); Clifton et al. (2012); Joyce et al. (2015); Koyama (2016); Nojiri et al. (2017); Capozziello et al. (2019); D’Agostino and Nunes (2019, 2022); D’Agostino and Bajardi (2025), ranging from higher-order curvature theories to extra-dimensional frameworks. Among these approaches, nonlocal theories of gravity Arkani-Hamed et al. (2002); Deser and Woodard (2007); Nojiri and Odintsov (2008); Hehl and Mashhoon (2009); Deffayet and Woodard (2009); Maggiore (2014) have emerged as a promising avenue to address fundamental inconsistencies of GR without introducing new physical fields or resorting to ad hoc corrections. Nonlocal modifications of GR have been motivated by various theoretical considerations, including their potential capability to address singularity problems Biswas et al. (2012); Buoninfante et al. (2019), as they offer a natural explanation for cosmic acceleration Maggiore and Mancarella (2014); Dirian et al. (2014); Capozziello et al. (2022); Capozziello and D’Agostino (2023), and reconcile gravity with QFT avoiding the traditional pitfalls of local higher-derivative theories Calcagni et al. (2007); Modesto (2012); Modesto and Rachwał (2017). Importantly, they provide novel phenomenological predictions that may be testable in future gravitational and cosmological observations, offering a pathway toward a deeper understanding of gravitational physics beyond Einstein’s theory.
@@
The aim of this paper is to investigate the physical implications of the newly derived BH solutions through the study of their quasinormal modes (QNMs), which serve as a powerful tool for probing the gravitational dynamics in strong-field regimes and identifying possible deviations from GR Dreyer et al. (2004). QNMs describe the characteristic oscillations of a perturbed BH, offering crucial insights into its stability and dynamics Chandrasekhar (1984); Berti et al. (2009). Mathematically, QNMs are characterized by complex frequencies, with real parts corresponding to oscillation frequencies and the imaginary parts related to the dumping timescales. A particular subclass of these fluctuations, namely the gravitational ones, govern the GW signals emitted by compact objects during merger and accretion processes Kokkotas and Schmidt (1999); Konoplya and Zhidenko (2011), providing thus a promising avenue to uncover potential observational signatures of nonlocal gravity.
@@ II A class of Black hole solutions in nonlocal gravity
Here, the first-order deviations from the Schwarzschild metric are measured by the small perturbative parameter $0<\alpha\ll 1$, while $k>1$ is a real constant D’Agostino and De Falco (2025a) [^1] Following the approach of D’Agostino and De Falco (2025a), in this work we express all physical quantities up to first-order in $\alpha$.. It can be shown that the spacetime arising from Eqs. (7a) and (7b) is free from essential singularities (except in $r=0$) beyond the event horizon and fulfills the asymptotic flatness condition (see D’Agostino and De Falco (2025a) for more details).
@@
where $x_{1}$ and $v_{1}$ are arbitrary real constants, and $\mathcal{P}_{k}(r)$ is a polynomial of order $32k-$ 3 in the radial coordinate $r$. Finally, the distortion function reads D’Agostino and De Falco (2025a)
@@ III.3 Gravitational perturbations
$e^{\lambda}_{\hat{\gamma}}\nabla_{\lambda}H_{\mu\nu}e_{\hat{\alpha}}^{\mu}e_{\hat{\beta}}^{\nu}=\partial_{\hat{\gamma}}H_{\hat{\alpha}\hat{\beta}}-\eta^{\hat{\delta}\hat{\epsilon}}\left(\Lambda_{\hat{\delta}\hat{\alpha}\hat{\gamma}}H_{\hat{\epsilon}\hat{\beta}}+\Lambda_{\hat{\delta}\hat{\beta}\hat{\gamma}}H_{\hat{\alpha}\hat{\epsilon}}\right)\,,$ (33)
@@
Now, we must evaluate the above equation in the background. Therefore, using Eqs. (29) and (30) and considering the $(\hat{\theta}\hat{\varphi})$ and $(\hat{r}\hat{\varphi})$ components, we obtain
@@ IV.2 Padé approximants
At this stage, it is important to note that the scalar and gravitational effective potentials (see Eqs. (55) and (56)), along with their derivatives, may exhibit numerical instabilities when evaluated near $r=3$. Hence, to avoid divergencies and ensure stable values for the QNMs, we employ the method of the Padé approximants Baker and Graves-Morris (1996) to smooth the behavior of these quantities around the peak. This technique has been widely used in the study of QNMs of BHs to enhance the accuracy and convergence of perturbative calculations (see e.g., Matyjasek and Opala (2017); Konoplya et al. (2019a); Konoplya and Stashko (2024) and references therein). The Padé approximant is defined as the ratio of two polynomials of order $p$ and $q$:
@@
[…]
@@ V Results and discussion
where $N_{s}$ is the total number of $(n,\ell)$ modes considered in Table 1 for a given $s$ [^8] For example, for $s=0$, we have $N_{0}=6$ and $\overline{\delta\omega}_{0}=\frac{1}{6}(\delta\omega_{000}+\delta\omega_{010}+\delta\omega_{110}+\delta\omega_{020}+\delta\omega_{120}+\delta\omega_{220})$. We report the resulting outcomes in Table 2 for different values of the perturbation parameter $\alpha$ and power-law index $k$. As expected, the departures from the Schwarzschild solution progressively decrease with increasing $k$ and decreasing $\alpha$ (see also Fig. 2). Specifically, for $\alpha=0.1$ and $k=1$, we find maximum deviations of 13.3%, 11.5%, and 11.8% for scalar, electromagnetic, and gravitational perturbations, respectively. These values decrease to $\sim 4\%$ for $\alpha=0.05$, and to $\sim 0.7\%$ for $\alpha=0.01$.
@@ VI Summary and perspectives
[…]
@@ Appendix A Effective potentials
Figure 4: Padé approximations of the effective potential and its derivatives for scalar perturbations, where $n=1$, $\ell=1$, $\alpha=0.1$, and $k=2$.
Figure 5: Padé approximations of effective potential and its derivatives for gravitational perturbations, where $n=1$, $\ell=1$, $\alpha=0.1$, and $k=2$.
@@ Appendix B Padé approximations
To illustrate the accuracy of the Padé method, we display in Figs. 4 and 5 the Padé approximations of the effective potential (and its derivatives) around $r=3$ for the scalar and gravitational BH perturbations, respectively. We observe that the Padé method enables an analytical treatment the divergences around the peak radius, while ensuring a smooth match with the original solutions outside this region. Although the precise degree of approximation is difficult to assess due to the presence of the divergences, Padé approximants have been shown to possess optimal convergence properties Capozziello et al. (2020), guaranteeing the reliability of our final results.

### v2

@@ I Introduction
Attempts to address these challenges have led to the development of modified gravity scenarios Ferraro and Fiorini (2007); Sotiriou and Faraoni (2010); Clifton et al. (2012); Joyce et al. (2015); Koyama (2016); Nojiri et al. (2017); Capozziello et al. (2019); D’Agostino and Nunes (2019, 2022); D’Agostino and Bajardi (2025), ranging from higher-order curvature theories to extra-dimensional frameworks. Among these approaches, nonlocal theories of gravity Arkani-Hamed et al. (2002); Deser and Woodard (2007); Nojiri and Odintsov (2008); Hehl and Mashhoon (2009); Deffayet and Woodard (2009); Maggiore (2014); Bombacigno et al. (2025) have emerged as a promising avenue to address fundamental inconsistencies of GR without introducing new physical fields or resorting to ad hoc corrections. Nonlocal modifications of GR have been motivated by various theoretical considerations, including their potential capability to address singularity problems Biswas et al. (2012); Buoninfante et al. (2019), as they offer a natural explanation for cosmic acceleration Maggiore and Mancarella (2014); Dirian et al. (2014); Capozziello et al. (2022); Capozziello and D’Agostino (2023), and reconcile gravity with QFT avoiding the traditional pitfalls of local higher-derivative theories Calcagni et al. (2007); Modesto (2012); Modesto and Rachwał (2017); Boos and Carone (2023). Importantly, they provide novel phenomenological predictions that may be testable in future gravitational and cosmological observations, offering a pathway toward a deeper understanding of gravitational physics beyond Einstein’s theory.
@@
The aim of this paper is to investigate the physical implications of the newly derived BH solutions through the study of their quasinormal modes (QNMs), which serve as a powerful tool for probing the gravitational dynamics in strong-field regimes and identifying possible deviations from GR Dreyer et al. (2004); Lin et al. (2024); Koshelev et al. (2024). QNMs describe the characteristic oscillations of a perturbed BH, offering crucial insights into its stability and dynamics Chandrasekhar (1984); Berti et al. (2009). Mathematically, QNMs are characterized by complex frequencies, with real parts corresponding to oscillation frequencies and the imaginary parts related to the dumping timescales. A particular subclass of these fluctuations, namely the gravitational ones, govern the GW signals emitted by compact objects during merger and accretion processes Kokkotas and Schmidt (1999); Konoplya and Zhidenko (2011), providing thus a promising avenue to uncover potential observational signatures of nonlocal gravity.
@@ II A class of Black hole solutions in nonlocal gravity
Here, the first-order deviations from the Schwarzschild metric are measured by the small perturbative parameter $0<\alpha\ll 1$, while $k>1$ is a real constant D’Agostino and De Falco (2025a) [^1] Following the approach of D’Agostino and De Falco (2025a), in this work we express all physical quantities up to first-order in $\alpha$.. It can be shown that the spacetime arising from Eqs. (7a) and (7b) is free from essential singularities (except in $r=0$) beyond the event horizon and fulfills the asymptotic flatness condition (see D’Agostino and De Falco (2025a) for more details). In fact, the Ricci scalar reads
$R(r)=\frac{3\alpha r^{-k-\frac{5}{2}}}{(r-3)^{3}}\bigg{[}3\left(7k^{2}-13k+4\right)r^{\frac{3}{2}}+k(k-1)r^{\frac{7}{2}}$
$+4k(3-2k)r^{\frac{5}{2}}-\frac{4r^{k+\frac{3}{2}}}{3^{k-1}}+\frac{2r^{k+\frac{1}{2}}}{3^{k-2}}-18(k-1)^{2}r^{\frac{1}{2}}\bigg{]},$ (8)
admitting the finite limit
$\lim_{r\to 3}R(r)=\frac{\alpha k\left(5+6k-k^{2}\right)}{3^{k+3}}\,.$ (9)
@@
where $x_{1}$ and $v_{1}$ are arbitrary real constants, and $\mathcal{P}_{k}(r)$ is a polynomial of order 2k32k-3 in the radial coordinate $r$. Finally, the distortion function reads D’Agostino and De Falco (2025a)
@@ III.3 Gravitational perturbations
[…]
@@
[…]
@@ IV.2 Padé approximants
Although the effective potentials are regular and well-defined at $r=3$ (see Eqs. (59), (60) and (61)), the numerical evaluation of their higher-order derivatives– required by the WKB method– can be affected by instabilities in the vicinity of that point. It is important to stress, however, that this behavior does not originate from any physical or geometrical singularity in the underlying spacetime. These numerical instabilities are common in the WKB analyses and can be effectively mitigated using standard techniques, such as Padé approximants Baker and Graves-Morris (1996), which improve the convergence and stability of perturbative calculations without altering the physical content of the potential. This method has been widely adopted in the study of QNMs of BH (see, e.g., Matyjasek and Opala (2017); Konoplya et al. (2019a); Konoplya and Stashko (2024) and references therein), where it helps to avoid spurious divergencies and to ensure stable QNM values by smoothing the behavior of the involved quantities around the potential peak.
The Padé approximant is defined as the ratio of two polynomials of order $p$ and $q$:
@@
where $x\equiv r-3$. To determine the unknown coefficients, we compute the Taylor series of the potential at x=0x=0777The $(p,q)$ Padé approximant must be matched to a Taylor series of order $p+q$.:
@@ V Results and discussion
where $N_{s}$ is the total number of $(n,\ell)$ modes considered in Table 1 for a given ss. [^8] For example, for $s=0$, we have $N_{0}=6$ and $\overline{\delta\omega}_{0}=\frac{1}{6}(\delta\omega_{000}+\delta\omega_{010}+\delta\omega_{110}+\delta\omega_{020}+\delta\omega_{120}+\delta\omega_{220})$. We report the resulting outcomes in Table 2 for different values of the perturbation parameter $\alpha$ and power-law index $k$. As expected, the departures from the Schwarzschild solution progressively decrease with increasing $k$ and decreasing $\alpha$ (see also Fig. 2). Specifically, for $\alpha=0.1$ and $k=1$, we find maximum deviations of 13.3%, 11.5%, and 11.8% for scalar, electromagnetic, and gravitational perturbations, respectively. These values decrease to $\sim 4\%$ for $\alpha=0.05$, and to $\sim 0.7\%$ for $\alpha=0.01$.
@@ VI Summary and perspectives
To quantify the departures from GR, we calculated the relative differences in the QNM frequencies with respect to the Schwarzschild solution, separating the real and the imaginary parts. For $\alpha=0.1$, the relative deviations from the Schwarzschild BH were found to be $\lesssim(13.3\%,\,11.5\%,\,11.8\%)$ for scalar, electromagnetic, and gravitational perturbations, respectively. As theoretically expected, the relative deviations progressively reduce with increasing $k$ and decreasing $\alpha$, reaching approximately zero for $k\simeq 5$ across all types of perturbation, independently of the specific value of $\alpha$.
@@ Appendix A Effective potentials
It is worth to emphasize that none of the above potentials exhibits a singularity at $r=3$, as demonstrated by the following finite limits:
$\lim_{r\to 3}V^{(0)}$ $=\frac{1}{2\times 3^{5+k}}\Big{\{}4\times 3^{1+k}+18\ell(\ell+1)\left(3^{k}-3\alpha\right)+\alpha\big{[}k\left(11+(k-6)k\right)-24\big{]}\Big{\}}\,$, (59)
$\lim_{r\to 3}V^{(1)}$ $=\frac{\ell(\ell+1)\left(3^{k}-3\alpha\right)}{3^{3+k}}\,$, (60)
$\lim_{r\to 3}V^{(2)}$ $=\frac{1}{486}\left\{18\left(\ell^{2}+\ell-2\right)-\frac{\alpha}{3^{k}}\left[k^{3}+17k+54\ell(\ell+1)-72\right]\right\}\,.$ (61)
Figure 4: Padé approximants of the derivatives of the effective potential for scalar perturbations for $n=1$, $\ell=1$, $\alpha=0.1$ and $k=2$.
Figure 5: Padé approximants of the derivatives of the effective potential for gravitational perturbations for $n=1$, $\ell=1$, $\alpha=0.1$ and $k=2$.
@@ Appendix B Padé approximations
To illustrate the accuracy of the Padé method, we display in Figs. 4 and 5 the Padé approximations of the derivatives of the effective potential for scalar and gravitational BH perturbations around $r=3$, respectively. We observe that the Padé method provides an effective analytical treatment of the numerical instabilities arising near the peak of the effective potential’s derivatives, while ensuring a smooth and accurate match with the original expressions away from this region. Although the precise degree of approximation is difficult to assess in this context, the Padé technique is known for its optimal convergence and stability properties Capozziello et al. (2020), which support the reliability of our final results.